OpenStudy (anonymous):

m+9/25=1/2 what is m?

OpenStudy (anonymous):

=> (25m +9)/ 25 = 1/2 => 25m +9 = 25/2

OpenStudy (anonymous):

the choises are 7/50 and 8/23 ?

OpenStudy (anonymous):

7/50 :)